a home is to be built on a rectangular plot with a perimeter of 800 feet. If the length is a 20 feet less than 3 times the width, what are the dimensions of the rectangular plot?

Answer by checkley79(3341)

P=2L+2W
L=3W-20
800=2(3W-20)+2W
800=6W-40+2W
800+40=8W
8408W
W=840/8
W=105 ANS.
L=3*105-20
L=315-20
L=295 ANS.
PROOF:
800=2*295+2*105
800=590+210
800=800